OpenStudy (anonymous):

Hi! Please help... a woman drops a front door key to her husband from their apartment window several stories above the ground. the function h=-16t^2 +64 gives the height h of the key in feet, t seconds after she releases it. A/ how long does it take the key to reach the ground? B/ what are the reasonable domain and range for the function h?

OpenStudy (anonymous):

that is a typical physics question :)

OpenStudy (anonymous):

a. solve for -16t^2 +64 = 0 -16t^2 = -64 t^2 = 4 t = 2 seconds

OpenStudy (anonymous):

thank you.! and what would be part B?

OpenStudy (anonymous):

my error was that i stayed in √4 lol

OpenStudy (anonymous):

D: \([0,2]\) R: \([0,64]\)

OpenStudy (anonymous):

how would i be getting to that conclusion?

OpenStudy (anonymous):

The key hits the ground at 2 seconds. Therefore, we are only concerned with those 2 seconds. The key travels from 64 feet above the ground to the ground. (Set t=0 to find the initial height.) So we are only concerned with those 64 feet.

OpenStudy (anonymous):

Oh ok thanks !